OpenStudy (anonymous):

I need help with five math questions . Any help is appreciated. Which is equation is an identity? (1 point) A. 8-(5x+2)=-5x-6 B. 7z+10-z=8z-2(z-5) C. 8m-4=5m+8-m D. 6y+5=6y-5

OpenStudy (anonymous):

An identity is an equality that holds regardless of the values of its variables. If you take left part of b) 7z+10-z is 6z+10 if you take right part of b) 8z-2(z-5)=8z-2z+10=6z+10 So left part and right part of the equality will always have the same value for whatever value of x you put. This is an identity

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

I ended up getting the right answer

OpenStudy (anonymous):

Do u just want the answer ???

OpenStudy (anonymous):

I already got the right answer to this like a week ago , but I do need someone to check my anwsers on these other problems.

OpenStudy (anonymous):

OK ! Now look : The answer is B because we have : 7z+10-z=8z-2(z-5) 7z+10-z=6z+10 8z-2(z-5)=8z-2z+10 So its right :) Got it :)?

OpenStudy (anonymous):

cool thanks , yea I got it. Mind helping me with these inequalities?

OpenStudy (anonymous):

x-7>10 My answer: x>17

OpenStudy (anonymous):

is this correct?

OpenStudy (anonymous):

Look : If u change the place of the negatives if we have > will have < Or have < have > Got it ?:)

OpenStudy (anonymous):

change the place of the negatives?

OpenStudy (anonymous):

Look :  Got it ?:)

OpenStudy (anonymous):

yea I got it.

OpenStudy (anonymous):

I still need (the answers to some/ to make sure I have the right answers to some) other problems though

OpenStudy (anonymous):

OK! Ask them :)

OpenStudy (anonymous):

sorry its taking so long I'm trying to write them all

OpenStudy (anonymous):

OK :)

OpenStudy (anonymous):

No problem :)

OpenStudy (anonymous):

@jojobugs :Do u have any problem or question ???

OpenStudy (anonymous):

yes i'm almost done typing them , thanks

OpenStudy (anonymous):

OK !

OpenStudy (anonymous):

2. 3x(underlined caret facing the left) 21 3. 5a-2<18 4. 2t+8(underlined caret facing the left) -4(t+1) Matching match the inequality to its graph Graphs A. -1 and under B. 3 to 0 C. 1 to -2 D. -3 to -6 5. 4y+3(underlined caret facing the right)y+6 Is the answer graph a,b,c,or d? 6. -2y>2 Is the answer graph a,b,c,or d? 7. y/3 < -1 Is the answer graph a,b,c,or d? 8. 3y(underlined caret facing the right) 2y+3 Is the answer graph a,b,c, or d?

OpenStudy (anonymous):

And your answers ??

OpenStudy (anonymous):

2.x(caret underlined facing the left)7 3. a<4 4. t(caret underlined facing the right)6 5.c 6.a 7.d 8.b

OpenStudy (anonymous):

For 3 please remeber your 2th question :)...

OpenStudy (anonymous):

in problems 2 and 4 I actually got 2. x(caret underlined facing the RIGHT)7 4. t(caret underlined facing the LEFT)6

OpenStudy (anonymous):

Whats the meaning of (caret underlined facing the RIGHT) ????

OpenStudy (anonymous):

greater than or equal to

OpenStudy (anonymous):

OK ! Now we have : 2t+8>=-4(t+1) 2t+8>=-4t-4 -6t+8=<-4 -6t>=-12 __________________________ 3x>=21 x>=7 OK ?:)

OpenStudy (anonymous):

okay thanks

OpenStudy (anonymous):

And do u want to know about other ???

OpenStudy (anonymous):

yes I would like to know what all the correct answers are

OpenStudy (anonymous):

Your problem is in the questions we have - things .

OpenStudy (anonymous):

what things?

OpenStudy (anonymous):

Negative !(-)

OpenStudy (anonymous):

o well what does that have to do with the anwsers to the problems?

OpenStudy (anonymous):

well thanks anyway

OpenStudy (anonymous):

Oh !:( No :( Im realy sorry friend :( Your answer for your last question and 2,4 was right :( I was in a big fall :( Excuse me :(...

OpenStudy (anonymous):

@jojobugs